OpenStudy (mollynicole2015):

Solve for p. -p + 10 = -9p – 6

OpenStudy (alekos):

first add p to both sides

OpenStudy (anonymous):

use math papa p= -2 step by step and solve it −p+10=−9p−6 Step 1: Add 9p to both sides. −p+10+9p=−9p−6+9p 8p+10=−6 Step 2: Subtract 10 from both sides. 8p+10−10=−6−10 8p=−16 Step 3: Divide both sides by 8. 8p 8 = −16 8 p=−2

OpenStudy (just_one_last_goodbye):

Hey Molly here is the steps to your question with the answer hope it helps ^_^ http://i.imgur.com/KmYDJjh.png

OpenStudy (alekos):

come on guys, your not supposed to just give out the entire solution